add and subtract the following
1) (7a^4+8ab+b^2)+(-5a^4-8ab-4b^2)
=7a^4+8ab+b^2-5a^4-8ab-4b^2
=(7a^4-5a^4)+(8ab-8ab)+(b^2-4b^2)
=2a^4+0-3b^2
=2a^4-3b^2 Ans
2) (6r+2s-1)-(4r-7s+1)
=(6r-4r)+(2s+7s)-(1+1)
=2r+9s-2 Ans
3) (7a^2b+4ab)+(9a^2b-8ab)-(11a^2b+14ab)
=(7a^2b+9a^2b-11a^2b)+(4ab-8ab-14ab)
=5a^2b-18ab
4) (-5x^3+2y^3)-(15x^3-7x^2y+8xy^2-4y^3)
=(-5x^3-15x^3)
question 1 I got 12a^4-3b^2
qustion 2 i got 2r-5s
qustion 3 i got 16a^2b+10ab-11a^2
question 4 I got -20x^3-xy^2
are these correct?

Found 2 solutions by KnightOwlTutor, amitsinghbce:
Answer by KnightOwlTutor(293)

1) (7a^4+8ab+b^2)+(-5a^4-8ab-4b^2)
(7-5)a^4+(1-4)b^2=2a^4-3b^2 The 8ab cancels out. You treated the 7 and 5 as if they had the same sign.Be careful.
2) (6r+2s-1-4r+7s-1)
(6-4)r+(2+7)s-2=2r+9s-2
Remeber to multiply out the negative sign -(4r-7s+1)=-4r+7s-1

3) (7a^2b+4ab)+(9a^2b-8ab)-11a^2b-14ab
(7+9-11)a^2b+(4-8-14)ab=5a^2b-18ab-I don't see an a^2 term here

4) (-5x^3+2y^3)-15x^3+7x^2y-8xy^2+4y^3)
(-5-15)x^3+7x^2y-8xy^2+(2+4)y^3)
-20x^3+7x^2y-8xy^2+6y^3 I really don't know how you got this answer
One ther is x^2y and the other is xy^2 they are not equivalent terms. Hope this helps.

question 1 I got 12a^4-3b^2
qustion 2 i got 2r-5s
qustion 3 i got 16a^2b+10ab-11a^2
question 4 I got -20x^3-xy^2
are these correct?

Answer by amitsinghbce(1)

1) (7a^4+8ab+b^2)+(-5a^4-8ab-4b^2)
=7a^4+8ab+b^2-5a^4-8ab-4b^2
=(7a^4-5a^4)+(8ab-8ab)+(b^2-4b^2)
=2a^4+0-3b^2
=2a^4-3b^2 Ans
2) (6r+2s-1)-(4r-7s+1)
=(6r-4r)+(2s+7s)-(1+1)
=2r+9s-2 Ans
3) (7a^2b+4ab)+(9a^2b-8ab)-(11a^2b+14ab)
=(7a^2b+9a^2b-11a^2b)+(4ab-8ab-14ab)
=5a^2b-18ab
4) (-5x^3+2y^3)-(15x^3-7x^2y+8xy^2-4y^3)
=(-5x^3-15x^3)+(2y^3-4y^3)-7x^2y+8xy^2
=-20x^3-2y^3-7x^2y+8xy^2 Ans